[Brahms-l] Shift report 20040201 08:00-16:00

From: Apache <apache@bnl.gov>
Date: Sun Feb 01 2004 - 15:49:27 EST
------------------------------------------------------------
Date: Sun Feb  1 15:49:27 2004, Shift: 08:00-16:00
Supervisor: sl, with: 
Log book#15 pages 47 - 47

Shift Summary:
------------------------------------------------------------
1.beam the entire shift except for about 45min after we voted for dumping.
2.Finished the 1,5k tr6 events at the 6deg A1219 setting
6ook tr3 events at 40deg B2255
3.No problems, Fleming tuned the trfs high voltages.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Feb 1 15:50:24 2004

This archive was generated by hypermail 2.1.8 : Sun Feb 01 2004 - 15:50:40 EST